• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name lkaufman

    Log entry time 16:06:03 on April11,2006

    Entry number 170419

    Followups:

    keyword=Day Shift Summary

    Day Shift Summary
    -----------------------
    SL: Lisa Kaufman
    TO: Klaus Grimm
    
    ----------------------
    We began the shift in production mode on 3He.  Beam studies started around 9 
    am.  Various activities going on in the hall, in particular, shielding was 
    added around big bite, and the target pumping polarization direction was 
    changed.  Beam studies ended around 3 pm and beam is not yet back at the end of 
    the shift.
    ----------------------
    
    Began shift taking 8 uA production mode data on 3He.
    
    09:12 - ended run 3838, HAPPEX run, and Compton run because beam studies 
    started.
    
    09:34 - changed BB, ND, and PS thresholds, BB HV, and prescalers.  Started a 
    cosmics run.
    
    10:05 - Hall A in restricted access.
    
    11:12 - Alex takes CODA for a while to test the Neutron detector.
    
    11:40 - Aidan and Ameya are going to the hall.  We will still be in restricted 
    access, but we will be going to laser controlled access, so only appropriately 
    trained people will be allowed into the hall at that time.
    
    12:15 - Aidan begins sweep of Hall A in order to go to laser controlled access.
    
    12:37 - Hall in Laser Controlled Access!
    
    13:47 - Started sweep of Hall A in order to go to Controlled access.
    
    14:27 - Hall in Controlled Access.
    
    15:30 - Hall going to Beam Permit.
    
    15:47 - Set detector thresholds and prescale factors back to production values.
    
    15:48 - Hall A in beam permit.
    
    Run Summary
    ---------------
    Run#  Time         Events   Current   IWHP   Type        Happex Run# 
    3837  07:38-08:43    2M       8uA      IN   Production       7331
    3838  08:44-09:11   700K      8uA      IN   Production       7332
    3839  09:34-10:25   380       0uA      IN    Cosmics
    3840  10:26-10:34    2M       0uA      IN    Cosmics (Neutron)
    3841  10:37-10:40   20K       0uA      IN    Cosmics
    3842  10:41-10:50    2M       0uA      IN    Cosmics (Neutron)
    3843  10:52-11:02   80K       0uA      IN    Cosmics
    3844  11:03-11:11   1.7M      0uA      IN    Cosmics (Neutron)
    
    Various other runs were taken while Alexandre Camsonne was in the hall working 
    on the neutron detector.